### v2

# Robust Trajectory Tracking of Autonomous Surface Vehicle via Lie Algebraic Online MPC

Abstract
Autonomous surface vehicles (ASVs) are influenced by environmental disturbances such as wind and waves, making accurate trajectory tracking a persistent challenge in dynamic marine conditions. In this paper, we propose an efficient controller for trajectory tracking of marine vehicles under unknown disturbances by combining a convex error-state MPC on the Lie group augmented by an online learning module to compensate for these disturbances in real time. This design enables adaptive and robust tracking control while maintaining computational efficiency. Extensive evaluations in the Virtual RobotX (VRX) simulator, and real-world field experiments demonstrate that our method achieves superior tracking accuracy under various disturbance scenarios compared with existing approaches.
I Introduction
The autonomous surface vehicle (ASV) has played an increasingly important role in the application of marine robots, undertaking a wide range of tasks such as hydrographic surveying, environmental monitoring, and water-quality samplings [22]. For surface vehicles, the time-varying hydrodynamic coefficients and external disturbances pose an obstacle to the controller’s trajectory-tracking ability, especially for underactuated ASV [13]. For example, winds, waves, currents, and inaccurate dynamics modeling can all disturb accurate tracking trajectory [7], which can lead to undesired consequences such as grounding, collision risks, or potential structural damage to the hull, thereby degrading reliability and operational safety.
Classical controllers like PID have been applied on marine robots [13], however, their performance is constrained under strong nonlinearities and time-varying disturbances. In contrast, Model Predictive Control (MPC) has demonstrated effectiveness in trajectory tracking, particularly in more complex environments [5] [21]. However, the non-linear hydrodynamics characteristics mean that the computation load of real-time MPC is heavy, resulting in a long response time to control signals, and a low motion frequency [16]. Moreover, small computing units lack the capacity to handle such heavy workloads, while large ones consume excessive space and energy, making them impractical for a compact ASV.
Multiple approaches have been put forward to alleviate the computational burden. In [12], MPC with projection neural network performs efficient constrained optimization problems with parallel computational capability. In [1], adaptive MPC leverages multiple approximated linear models to simplify the calculation. Another approach is to achieve geometric control [8], based on a Lie group framework, to exploit symmetry in the problem. In this concept, Lie Algebraic MPC (Lie-MPC) is proposed [19], and applied on ASV [10]. Lie-MPC reduces computation time, making online learning for non-linear hydrodynamics applicable to ASV.
Figure 1: Our method employs Fourier features to estimate the residual in an online manner and can be implemented efficiently on a single-core CPU. In contrast, prior online approaches rely on neural networks that require pre-training on large datasets.
In parallel, several online learning methods have been proposed. In [20], a neural network–based adaptive module is incorporated into a model‐based controller to compensate for large wave forces. In [14],the Deep Neural Network (DNN) is first trained offline to learn the model dynamics, then it is updated online to fit unknown disturbances. Similarly, in [15], a DNN is used to learn the system dynamics, and an Extended State Observer (ESO) measures error and makes up for residual dynamics estimation in the kinematic control law.
However, these neural‐network approaches generally require GPU resources for offline pre‐training on large datasets, which increases computational cost. Moreover, like in [15], the online-learning pipeline is used to directly change the system dynamics, making it difficult to generalize across different system setups. In [23], all disturbances are unified into a single locally Lipschitz term within the MPC dynamics and estimated through feature-based modeling. While this approach has been demonstrated on aerial vehicles and shown to reduce computational burden, its applicability to ASVs and its interaction with a Lie-MPC controller remain unexplored.
To mitigate the research gap, the main contributions of this work are as follows:
  1. 1.
    A real-time online learning framework for MPC on ASVs. We propose a novel integration of Fourier-based online learning with convex Lie algebraic MPC that enables real-time disturbance compensation using only a single CPU core.
  2. 2.
    Bi-level feature extraction for efficient disturbance representation. We develop a bi-level Fourier-based feature extraction method that identifies and retains the dominant components of disturbance dynamics.
  3. 3.
    Validation across simulation and real-world domains. The proposed controller is evaluated in a physics-based simulator and in real-world experiments.
The remainder of this paper is organized as follows. Section II provides the mathematical preliminaries. Section III presents the details of our proposed method. Experiments are presented in Section IV. Finally, Section V concludes the paper and discusses future studies.
II Preliminaries and Mathematical Formulation
This section provides a brief overview of the necessary mathematical background used in the developed approach.
Figure 2: This figure illustrates the overall pipeline. At each control cycle, a residual term is first computed by comparing the measured state derivative with the nominal dynamics model. A sliding buffer containing the most recent residual samples serves as the input to the Fourier-based online estimator. Then, the estimator updates the weight matrix to approximate this residual-history buffer and predicts a full horizon of future residual terms. These predicted residuals are then injected into the Lie-MPC, allowing the controller to account for upcoming disturbances over the entire planning horizon. In this manner, the proposed framework achieves real-time disturbance compensation without any weight pre-training and operates efficiently on resource-limited ASV platforms.
II-A Dynamics on Lie Group
This section briefly introduces the rigid body dynamics on the Lie group $\mathrm{SE}(3)$. For more backgrounds on matrix Lie groups, refer to [9, 18, 3].
Let $\mathcal{G}$ be an $n$-dimensional matrix Lie group with its associated Lie algebra $\mathfrak{g}$. We define the following isomorphism
$(\cdot)^{\wedge}:\mathbb{R}^{n}\rightarrow\mathfrak{g},$ (1)
that maps an element in the vector space $\mathbb{R}^{n}$ to the tangent space of the matrix Lie group at the identity. Its inverse mapping is denoted as $(\cdot)^{\vee}:\mathfrak{g}\rightarrow\mathbb{R}^{n}$. Then, for any $\phi\in\mathbb{R}^{n}$, we define the exponential map as
$\exp(\cdot):\mathbb{R}^{n}\rightarrow\mathcal{G},\ \ \exp(\phi)=\operatorname{exp_{m}}({\phi}^{\wedge}),$ (2)
where $\operatorname{exp_{m}}(\cdot)$ denotes the matrix exponential.
To describe the 3D rigid body motion of the vehicle, we represent the rotation by matrix $R\in\mathrm{SO}(3)$ and the position vector by $p\in\mathbb{R}^{3}$ and have the full state as:
$X=\begin{bmatrix}R&p\\ 0&1\end{bmatrix}\in\mathrm{SE}(3).$ (3)
The tangent space at the identity, $\mathfrak{se}(3)$, consists of twist $\xi\in\mathbb{R}^{6}$, which is defined as the concatenation of the angular velocity $\omega$ and the linear velocity $v$ in the body frame:
$\xi=\begin{bmatrix}\omega\\ v\end{bmatrix},\quad\xi^{\wedge}=\begin{bmatrix}\omega^{\wedge}&v\\ 0&0\end{bmatrix}\in\mathfrak{se}(3).$ (4)
The adjoint map of $\mathrm{SE}(3)$ has the matrix representation:
$\mathrm{Ad}_{X}=\begin{bmatrix}R&0\\ p^{\wedge}R&R\end{bmatrix},\qquad\mathrm{ad}_{\xi}=\begin{bmatrix}\omega^{\wedge}&0\\ v^{\wedge}&\omega^{\wedge}\end{bmatrix}$, (5)
where $\mathrm{Ad}_{X}$ describes the change of frame of Lie algebra, and $\mathrm{ad}_{\xi}$ is defined by $\mathrm{ad}_{\xi}\eta=[\xi,\eta]$ with $\cdot$$\cdot$] the Lie bracket.
II-B Dynamics of ASV in Fossen’s model
We will use a 6-DOF motion equations based on Fossen’s model [6] to describe the ASV, with the the prescribed Lie group’s ordering of twist vector $\xi=[\omega,v]$. The dynamics are represented in the body frame following the convention of Euler Poincare equation for single rigid body:
$M_{RB}\dot{\xi}$ $+C_{RB}(\xi)\xi+M_{AM}\dot{\xi}_{r}+C_{AM}(\xi_{r})\xi_{r}$
$+D(\xi_{r})\xi_{r}+g(\eta)=\tau_{c}+\tau_{disturb}^{body}$ (6)
with $M_{(\cdot)}$ the mass matrix, $C_{(\cdot)}$ the Coriolis matrix. The subscript $RB$ refers to the rigid-body contribution, while $AM$ means “added mass”, which represents the inertia of fluid accelerated by the vehicle motions. $D(\xi_{r})$ is the hydrodynamic damping (drag) matrix. $g(\eta)$ is the hydrostatic restoring force and moment vector, and here we assume the vehicle to be neutrally buoyant, so the hydrostatic term $g(\eta)$ becomes negligible. $\xi_{r}=\xi-\xi_{c}$ is the relative twist with respect to the ocean current $\xi_{c}$. $\tau_{c}$ is the six-dimensional control input from the propulsion system, $\tau_{disturb}^{body}$ is the six-dimensional external disturbance transformed in the body frame.
The mass matrix corresponding to $\xi$ is given by
$M=\begin{bmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{bmatrix},\quad M_{ij}\in\mathbb{R}^{3\times 3},$ (7)
there the submatrices are defined as $M_{11}$ is a positive definite inertia matrix in body frame, $M_{12}=M_{21}^{\top}=-m\,r_{g}^{\wedge}$, and $M_{22}=mI_{3}$, where $r_{g}=[x_{g},y_{g},z_{g}]^{\mathsf{T}}$ denotes the position vector of the center of gravity expressed in the body-fixed frame. $(\cdot)^{\wedge}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ denotes the wedge operator that maps a vector to its corresponding skew-symmetric matrix. The Coriolis matrix is given by
$C(\xi)=\begin{bmatrix}-\left(M_{11}\omega+M_{12}v\right)^{\wedge}&-\left(M_{21}\omega+M_{22}v\right)^{\wedge}\\ -\left(M_{11}\omega+M_{12}v\right)^{\wedge}&0\end{bmatrix}$ (8)
which is a skew-symmetric that represents the “inertial forces” in the vehicle dynamics. The damping matrix is given by
$D(\xi_{r})$ $=-\,\mathrm{diag}\!\left(\begin{bmatrix}K_{p},\;M_{q},\;N_{r},\;X_{u},\;Y_{v},\;Z_{w}\end{bmatrix}^{\mathsf{T}}\right)$ (9)
$\quad-\,\mathrm{diag}\!\left(\begin{bmatrix}K_{|p|p}|p|,\;M_{|q|q}|q|,\;N_{|r|r}|r|,\\[6.0pt] X_{|u|u}|u|,\;Y_{|v|v}|v|,\;Z_{|w|w}|w|\end{bmatrix}^{\mathsf{T}}\right)$
The damping terms model hydrodynamic effects such as potential damping, skin friction, wave drift damping, vortex shedding, and lifting forces. Depending on the surge velocity, either linear or nonlinear damping terms dominate. For non-coupled motion, a diagonal damping structure can be assumed, simplifying the dynamics.
Combining (5) with (6), have the full dynamics
$M\dot{\xi}=\mathrm{ad}^{\mathsf{T}}_{\xi}M\xi+F=-C(\xi)\xi-D(\xi)\xi+\tau,$ (10)
$\begin{bmatrix}\dot{R}&\dot{p}\\ 0&0\end{bmatrix}=\begin{bmatrix}R&p\\ 0&1\end{bmatrix}\begin{bmatrix}\omega^{\wedge}&v\\ 0&0\end{bmatrix}$. (11)
with $F\in\mathfrak{se}^{*}(3)$ the external force in the body frame.
III Methods
In this section, we
III-A Error-State Convex MPC
We develop a convex error-state MPC by linearizing the tracking error dynamics in the Lie algebra and learn the residual dynamics to compensate for the modeling error.
Let the desired trajectory be $X_{d,t}\in\mathrm{SE}(3)$, and define the left-invariant error [2] on matrix Lie group as:
$\Psi=X_{d,t}^{-1}X_{t},\quad\dot{\Psi}=\Psi(\xi_{t}-Ad_{\Psi_{t}^{-1}}\xi_{d,t})^{\wedge}.$ (12)
Using the first-order approximation $\Psi_{t}\approx I+\psi_{t}^{\wedge}$, we obtain the linearized error dynamics in Lie algebra:
$\dot{\psi}_{t}=-ad_{\xi_{d,t}}\psi_{t}+\xi_{t}-\xi_{d,t}.$ (13)
Linearizing the hydrodynamic model around $\bar{\xi}=\xi_{d,t}$ yields:
$M\dot{\xi}=H_{t}\xi+b_{t}+\delta_{t}^{d}+\tau,$ (14)
where $H_{t}$ and $b_{t}$ are the Jacobian and residual terms derived from the first-order expansion of $C(\xi)$ and $D(\xi)$, while $\delta_{t}^{d}$ is derived from external hydrodynamic disturbance.
Define the system state $x_{t}=[\psi_{t},\xi_{t}]^{\mathsf{T}}$, and the linearized hydrodynamic model becomes:
$\dot{x}_{t}=A_{t}x_{t}+B_{t}\tau+h_{t}+\delta_{t},$ (15)
with each component defined by:
$A_{t}$ $=\begin{bmatrix}-ad_{\xi_{d,t}}&I\\[4.0pt] 0&H_{t}\end{bmatrix},\quad B_{t}=\begin{bmatrix}0\\[4.0pt] M^{-1}\end{bmatrix}$, (16)
$h_{t}$ $=\begin{bmatrix}-\xi_{d,t}\\[4.0pt] b_{t}\end{bmatrix},\quad\delta_{t}=\begin{bmatrix}\delta_{t}^{e}\\[4.0pt] \delta_{t}^{d}\end{bmatrix}$.
Here, $\delta_{t}^{e}$ represents the higher-order linearization residuals in the error-state dynamics, capturing higher-order terms neglected in the first-order approximation around $\xi_{d,t}$, $\delta_{t}^{d}$ denotes the external disturbances and unmodeled hydrodynamic effects.
For cost function, we first define the tracking output:
$y_{t}=\begin{bmatrix}\psi_{t}\\ \dot{\psi}_{t}\end{bmatrix}=G_{t}x_{t}-d_{t},$ (17)
where
$G_{t}=\begin{bmatrix}I&0\\ -ad_{\xi_{d,t}}&0\end{bmatrix},\quad d_{t}=\begin{bmatrix}0\\ \xi_{d,t}\end{bmatrix}$. (18)
Then the cost function is formulated as:
$J=y_{N}^{\mathsf{T}}Py_{N}+\sum_{k=1}^{N-1}(y_{k}^{\mathsf{T}}Qy_{k}+\tau_{k}^{\mathsf{T}}R\tau_{k}),$ (19)
with $P$, $Q$, and $R$ the positive definite weighting matrices.
After discretization with step $\Delta t$, the MPC optimization problem becomes:
$\min_{\tau_{k}}$ $y_{N}^{\mathsf{T}}Py_{N}+\sum_{k=1}^{N-1}(y_{k}^{\mathsf{T}}Qy_{k}+\tau_{k}^{\mathsf{T}}R\tau_{k})$ (20)
s.t. $x_{k+1}=A_{k}x_{k}+B_{k}\tau_{k}+h_{k}+\delta_{k},$
$x_{0}=x_{\text{init}},\quad\tau_{k}\in\mathcal{T}_{k},$
where $A_{k}=I+A_{t_{k}}\Delta t$, $B_{k}=B_{t_{k}}\Delta t$, $h_{k}=h_{t_{k}}\Delta t$, and and $\delta_{k}=\delta_{t_{k}}\Delta t$. The resulting QP can be efficiently solved using the OSQP solver.
Algorithm 1 Fourier Online MPC
1:Total timesteps $T$; MPC prediction horizon NN; feature map $\Phi(\cdot)$; step size $\eta$; regularizer $\lambda$
[…]
5:for $t=1,\ldots,T$ do
6:  $\{\hat{\delta}_{t+k|t}\}_{k=1}^{H}\;\leftarrow\;\{\hat{C}_{t}^{\top}\Phi(Z_{t+k|t})\}_{k=1}^{H}\triangleright$ Predict Residual
7:  $\tau_{t}\;\leftarrow\;\mathrm{MPC}\!\big(x_{t},\,\{\hat{\delta}_{t+k}\}_{k=1}^{H}\big)$ $\triangleright$ Solve (20)
8:  $\delta_{t}=\dot{x}_{t}-(A_{t}x_{t}+B_{t}\tau_{t}+h_{t})$ $\triangleright$ Estimate residual
[…]
10:  $\hat{C}_{t+1}\leftarrow\hat{C}_{t}-\eta\nabla_{\hat{C}_{t}}\ell_{t}$ $\triangleright$ Update weights
11:end for
III-B Fourier-based Online Learning
At each time step $t$, we estimate the residual term $\delta_{t}$ by comparing the measured and nominal state derivatives:
$\delta_{t}=\dot{x}_{t}-(A_{t}x_{t}+B_{t}\tau_{t}+h_{t}),$ (21)
where $\dot{x}_{t}=(x_{t+1}-x_{t})/\Delta t$ is obtained from discrete-time measurements using GPS and IMU fusion[17], and $(A_{t}x_{t}+B_{t}\tau_{t}+h_{t})$ is the nominal model from (15). The residual $\delta_{t}=[\delta_{t}^{e},\delta_{t}^{d}]^{\top}$ captures both internal linearization errors and external disturbances.
We approximate the residual by a Fourier basis evaluated on an input that concatenates the state $x_{t}$ and time $t$:
$Z_{t}=(x_{t},t),\qquad\hat{\delta}_{t}={C}^{\top}\,\Phi(Z_{t}),$ (22)
where $\Phi(\cdot):\mathbb{R}^{12}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the Fourier feature map that maps the time-dependent robot state to a $d$-dimensional feature. The coefficient matrix C{C} is updated online using samples stored in a sliding buffer $\mathcal{B}$.
We use the most recent samples to compute a buffered loss with a temporal smoothness regularizer:
$\ell_{t}\;=\;\frac{1}{N}\!\sum_{j=t-N+1}^{t}\Big(\,\|\delta_{j}-\hat{\delta}_{j}\|^{2}+\lambda\,\|\hat{\delta}_{j}-\hat{\delta}_{j-1}\|^{2}\Big),$ (23)
We set $\hat{\delta}_{0}\!=0$ for the smoothness term when $j=1$. The optimization problem is formulated as
$\hat{C}=\arg\min_{C}\quad\ell_{t}$ (24)
III-C Offline Bi-level Feature Extraction
Inspired by the Fourier-based nature of wave and wind disturbances [11], the features we use are all Fourier-based. However, directly using a large number of random Fourier features for disturbance estimation causes many of them fail to capture the dominant frequency components and spatial correlations of the residual term. To address this issue, we perform feature extraction to identify and retain components that contribute significantly to the residual representation. This reduces both the computational cost and the risk of overfitting while preserving model accuracy.
Each feature is represented as a sine–cosine pair obtained from a single variable $z_{t,i}$ with a trainable frequency $f_{j}$:
$\Phi(Z_{t})=\big[\sin(f_{j}\,z_{t,i}),\;\cos(f_{j}\,z_{t,i})\big]_{i=1,\dots,d;\,j=1,\dots,m}$ (25)
where $d$ is the state dimension (12-dimensional state and 1-dimensional time) and $m$ is the number of selected frequencies.
Algorithm 2 Offline Bi-level Feature Extraction
1:Dataset $\{(Z_{t},h_{t})\}_{t=1}^{T}$; initial amplitudes $\hat{C}$; initial frequency set $\{f_{j}\}_{j=1}^{M}$; feature map $\Phi(\cdot)$; inner steps $n_{\mathrm{inner}}$; epochs NN; learning rate $\eta$; regularization $\lambda$
2:Output: Optimized Fourier feature map $\Phi(\cdot)$
3:for $\text{epoch}=1$ to NN do
4:  Freeze $\{f_{j}\}$; unfreeze $\hat{C} \triangleright$ Inner loop
5:  for $k=1$ to $n_{\mathrm{inner}}$ do
6:   for all $(Z_{t},h_{t})$ in dataset do
7:     $\hat{h}_{t}\leftarrow\hat{C}^{\top}\Phi(Z_{t})$
8:     $L_{\mathrm{inner}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}+\lambda\|\hat{C}\|_{1}$ $\triangleright$ Compute loss
9:     $\hat{C}\leftarrow\hat{C}-\eta\nabla_{\hat{C}}L_{\mathrm{inner}}$ $\triangleright$ Update weights
10:   end for
11:  end for
12:  Freeze $\hat{C}$; unfreeze $\{f_{j}\}$ $\triangleright$ Outer loop
13:  for all $(Z_{t},h_{t})$ in dataset do
14:   $\hat{h}_{t}\leftarrow\hat{C}^{\top}\Phi(Z_{t})$
15:   $L_{\mathrm{outer}}\leftarrow\|\hat{h}_{t}-h_{t}\|_{1}$ $\triangleright$ Compute loss
16:   $\{f_{j}\}\leftarrow\{f_{j}\}-\eta\nabla_{f_{j}}L_{\mathrm{outer}}$ $\triangleright$ Update weights
17:  end for
18:end for
[…]
Let $\{(Z_{t},h_{t})\}_{t=1}^{T}$ denote the dataset, where $T$ is the total number of samples.We use a bi-level optimization to extract features, which alternates between amplitude and frequency updates. The problem is formulated as
$\min_{\{f_{j}\}}\;\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\Big(h_{t},\;\hat{C}(f)\,\Phi(Z_{t})\Big).$ (26)
where
$\hat{C}(f)\;=\;\arg\min_{C}\frac{1}{T}\sum_{t=1}^{\mathsf{T}}\ell\!\big(h_{t},\,C\,\Phi(Z_{t})\big)+\lambda\|C\|_{1}.$ (27)
with $\ell(\cdot,\cdot)$ denoting the L2 prediction loss. The inner optimization loop optimizes the amplitude parameters $C$ with an $\ell_{1}$ regularization that promotes sparsity, while the outer optimization loop adjusts the frequencies $\{f_{j}\}$ to minimize the overall prediction error.
The inner optimization is a convex problem with respect to $C$, which guarantees a unique global optimum for fixed frequencies. We solve the overall bilevel objective using alternating minimization: fix $\{f_{j}\}$ and update $C$ (inner loop), then fix $C$ and update $\{f_{j}\}$ (outer loop). This alternating scheme adaptively learns both the most informative frequencies and their corresponding amplitudes. Even in the presence of Gaussian noise in the disturbance data, the convexity of the inner layer ensures convergence to a stable global minimum. The implementation is detailed in Algorithm 2.
IV Experiments
IV-A Experiment Setup
IV-A1 Repetition Protocol
This section involves tests under two scenarios: (i) the Virtual RobotX (VRX) simulator, and (ii) a real-world river experiment under wind. Tests in scenarios (i) are each repeated 10 times, with an initial position offset $\Delta\mathbf{p}_{0}=[\,1.0\sqrt{u}\cos\theta,\;1.0\sqrt{u}\sin\theta,\;0\,]^{\top}\mathrm{m}$, where $\theta\sim\mathcal{U}(0,2\pi)$ and $u\sim\mathcal{U}(0,1)$.
IV-A2 Controller Tuning
To reduce the coupling between observer and controller, we first approximate the dynamics with an LQR approximation. Given weighting matrices $Q$, $R$, we compute $P$ from the discrete-time algebraic Riccati equation (DARE)
$P=A^{\top}PA=A^{\top}PB(R+B^{\top}PB)^{-1}B^{\top}PA+Q,$ (28)
and set the stabilizing gain
$K=(R+B^{\top}PB)^{-1}B^{\top}PA.$ (29)
We tune weighting matrices $Q$ and $R$ based on eigenvalues of ABKA-BK, avoiding complex conjugate pairs and ensuring the norm remains well below unity.
IV-A3 Baselines
We compare the performance of our method with those baseline methods: (1) Nominal Lie-MPC: the Lie-MPC framework without online learning [10]. (2) L1 Adaptive-MPC: the same MPC framework with an online L1-style disturbance–estimation module to calculate compensation based on residual dynamics as well. [4] (3) PID: A two-layer PID controller with an outer loop that computes the desired twist from pose errors, and an inner loop that tracks these references using two independent PID controllers:
$v_{x}^{\mathrm{ref}},\,\omega_{z}^{\mathrm{ref}}=f_{\mathrm{outer}}(e_{x},e_{y},e_{\psi}),\\$ (30)
$u=f_{\mathrm{inner}}\!\left(\mathrm{PID}_{v}(v_{x}^{\mathrm{ref}}-v_{x}),\,\mathrm{PID}_{\omega}(\omega_{z}^{\mathrm{ref}}-\omega_{z})\right),$ (31)
where $u$ denotes the differential thrust command.
IV-A4 Features for disturbance estimation
After collecting disturbance and corresponding state variable in 10 rounds of simulation run in the Fossen model. We employ Algorithm 2 to extract 30 features, which are then used in multiple scenarios under various disturbances. The extracted features exhibit strong generalization capability.
IV-A5 Performance metric
We use the Root Mean Square Error (RMSE) as the performance metric for planar tracking. Since a random initial position error is introduced in each test, the RMSE is evaluated only after the ASV reaches a steady state to exclude the relatively large error at the beginning of each run. For instance, the RMSE is calculated over the second half of the path. For the selected segment $\mathcal{K}$, the trajectory-tracking RMSE for each run is
$\mathrm{RMSE}=\sqrt{\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\big\|[x_{k},\,y_{k}]^{\top}-[x_{k}^{\mathrm{ref}},\,y_{k}^{\mathrm{ref}}]^{\top}\big\|_{2}}.$
The values in the table are the mean and standard deviation of $\mathrm{RMSE}$ across repeated trials.
IV-A6 Implementation details
The experiments are conducted using two reference trajectories: (i) a zigzag path defined by a time-varying yaw rate $\omega_{z}(t)=0.1\cos(t/100)$ with constant surge speed $v_{x}=0.5\,\mathrm{m/s}$, and (ii) a lawn-mower coverage pattern characterized by repeated straight segments ($v_{x}=0.5\,\mathrm{m/s}$) and 90° turns executed at a turning rate $\omega_{z}=0.3\,\mathrm{rad/s}$. Both trajectories have a total duration of $128\,\mathrm{s}$.
To incorporate the learned disturbances into the global MPC model, the estimated body-frame disturbance wrench $\tau_{t}^{\text{body}}$ is transformed into the world frame via
$\tau_{t}^{\text{world}}=\begin{bmatrix}R(X_{t})&0\\ 0&R(X_{t})\end{bmatrix}\tau_{t}^{\text{body}},$ (32)
where $R(X_{t})$ is the rotation matrix corresponding to the ASV orientation. This operation assumes spatially uniform disturbances acting at the center of mass.
The algorithm configuration used across all experiments is summarized in Table I.
Table I: Algorithm Configuration
Control Rate Control Horizon NN step size $\eta$ Regularizer $\lambda$
50Hz 30 $10^{-4}$ $10^{-4}$
Figure 3: The VRX environment.
Figure 4: Tracking performance of baselines and our method under different wind conditions in the VRX simulator.
Figure 5: Tracking performance of baselines and our method under different wind conditions in the VRX simulator.
IV-B Simulations in physics-based Virtual RobotX simulator
IV-B1 System Setup
We evaluate algorithm performance using the Virtual RobotX (VRX) simulator, which provides a physics-based marine environment with realistic hydrodynamic and wind effects, as shown in Fig. 3. In simulations, the wave model and random wind seed are disabled for consistency, while the wind gain is fixed at 1 under the constant-speed option. A URDF model of the real-world ASV is imported into the simulator; the vessel has a mass of 28 kg and is actuated by two fixed parallel propellers with thrust limits of [26,50][-26,50] N.
Table II: Trajectory-tracking RMSE for zigzag trajectory in VRX simulator.
Controller 0.0 m/s 1.0 m/s 2.0 m/s 3.0 m/s
PID $2.901\pm 0.068$ $3.115\pm 0.044$ $3.932\pm 0.175$ $8.798\pm 0.334$
MPC $\mathbf{0.093\pm 0.002}$ $0.193\pm 0.051$ $0.809\pm 0.013$ $1.441\pm 0.710$
L1 $098\pm 0.052 1.1$$165\pm 0.163$ $1.463\pm 0.275$ $2.319\pm 0.215$
Ours $0.103\pm 0.002$ $\mathbf{0.135\pm 0.003}$ $\mathbf{0.406\pm 0.008}$ $\mathbf{0.830\pm 0.241}$
Table III: Trajectory-tracking RMSE for lawnmower trajectory in VRX simulator.
Controller 0.0 m/s 1.0 m/s 2.0 m/s 3.0 m/s
PID $2.131\pm 0.001$ $2.137\pm 0.001$ $159\pm 0.002 2.2$$209\pm 0.002$
MPC $2\pm 0.001 0.022±0.0$$2\pm 0.001$ $0.046\pm 0.001$ $0.131\pm 0.000$
L1 $0.079\pm 0.015$ $107\pm 0.052 0.0$$75\pm 0.032 0.0$$075\pm 0.015$
Ours $\mathbf{0.016\pm 0.003}$ $\mathbf{0.017\pm 0.003}$ $\mathbf{0.029\pm 0.002}$ $\mathbf{0.069\pm 0.001}$
IV-B2 Results
As shown in Table II, Table III, Fig. 4, and Fig. 5, the proposed Online-MPC consistently achieves the lowest tracking error across all evaluated wind conditions and trajectories. For both the zigzag and lawnmower trajectories, Online-MPC outperforms the nominal MPC, with the performance gap widening as wind speed increases, demonstrating improved robustness and disturbance-rejection capability under stronger environmental disturbances.
The PID controller exhibits the highest tracking error overall. Its performance degrades substantially with increasing wind speed, particularly in the zigzag trajectory, where the RMSE rises from approximately 2.9 at 0 m/s to nearly 8.8 at 3 m/s. Even in the simpler lawnmower trajectory, PID maintains RMSE above 2.1 across all wind conditions, showing limited adaptability to disturbances. Similarly, the L1 adaptive MPC performs worse than nominal-MPC, possibly due to over-estimation under low disturbance. This pattern can also be observed in the lawnmower trajectory, where L1 performs poorly under low wind disturbance but outperforms nominal-MPC in 3 m/s wind.
IV-C Real-world Field Test
IV-C1 System Setup
As shown in Fig. 6, the ASV used is a modified Blue Robotics BlueBoat. We leverage an Nvidia Jetson Orin AGX to handle on-board computation. The vessel, together with a sensor suite, has a mass of 28 kg and is actuated by two fixed parallel propellers with thrust limits of [26,50][-26,50] N.
Field experiments were conducted in the same environment multiple times. We launch the boat in three different directions, and separately test the performance of the proposed method and baseline Lie-MPC. External disturbances in the experiments include wind (approximately 1 m/s) and minor drag from the safety rope. These effects, together with internal linearization errors, provide a realistic evaluation of controller robustness.
Figure 6: Real-world field test setup.
Figure 7: The trajectories in the field test.
IV-C2 Results
As shown in Table IV, Fig. 7, the proposed Online-MPC consistently outperforms the nominal MPC in zigzag trajectories. It achieves lower tracking errors and maintains stable operation under natural wind and current disturbances, demonstrating improved robustness and disturbance rejection capability in real-world conditions.
Table IV: Mean RMSE of trajectory tracking in field test.
Controller RMSE (m)
MPC $1.17\pm 0.85$
Ours $\mathbf{0.76\pm 0.56}$
V Conclusion
We presented an Online Learning-based Model Predictive Control (Online-MPC) framework for trajectory tracking under unknown disturbances. By integrating online residual compensation into MPC, the controller compensates for unmodeled dynamics such as wind and hydrodynamic variations in real time.
Both simulation and field experiments demonstrate that Online-MPC consistently outperforms the baselines, achieving lower tracking errors and greater robustness across diverse operating conditions. The results highlight the method’s effectiveness and practicality as a low-cost method for autonomous surface vessels operating in uncertain and dynamic environments.
Further study may focus on deploying this algorithm on other ASVs in an environment with obvious waves or evaluating the algorithm on a more task-based measurement for common ASV missions.